### v1

# S1-Bench: A Simple Benchmark for Evaluating System 1 Thinking Capability of Large Reasoning Models

Abstract
We introduce S1-Bench, a novel benchmark designed to evaluate Large Reasoning Models’ (LRMs) performance on simple tasks that favor intuitive system 1 thinking rather than deliberative system 2 reasoning. While LRMs have achieved significant breakthroughs in complex reasoning tasks through explicit chains of thought, their reliance on deep analytical thinking may limit their system 1 thinking capabilities. Moreover, a lack of benchmark currently exists to evaluate LRMs’ performance in tasks that require such capabilities. To fill this gap, S1-Bench presents a set of simple, diverse, and naturally clear questions across multiple domains and languages, specifically designed to assess LRMs’ performance in such tasks. Our comprehensive evaluation of 22 LRMs reveals significant lower efficiency tendencies, with outputs averaging 15.5 times longer than those of traditional small LLMs. Additionally, LRMs often identify correct answers early but continue unnecessary deliberation, with some models even producing numerous errors. These findings highlight the rigid reasoning patterns of current LRMs and underscore the substantial development needed to achieve balanced dual-system thinking capabilities that can adapt appropriately to task complexity [^1] The code and benchmark can be found in https://github.com/WYRipple/S1_Bench..
[…]
tcb@breakable
S1-Bench: A Simple Benchmark for Evaluating System 1 Thinking Capability of Large Reasoning Models
Wenyuan Zhangthanks:  denotes equal contribution. denotes corresponding author., Shuaiyi Nie, Xinghua Zhang, Zefeng Zhang, Tingwen Liu Institute of Information Engineering, Chinese Academy of Sciences School of Cyber Security, University of Chinese Academy of Sciences {zhangwenyuan,nieshuaiyi,liutingwen}@iie.ac.cn
Figure 1: LRMs exhibit under-accuracy and overthinking on simple problems. Shapes represent organizations, colors represent base model families, with darker colors indicating larger models, and connecting lines represent the relationships between model families and training.
1 Introduction
“Simplicity is the ultimate sophistication.”
— Leonardo da Vinci
Recent advances in Large Reasoning Models (LRMs), notably OpenAI’s o1/o3 OpenAI (2024) and the DeepSeek-R1 Guo et al. (2025) series, have propelled the development of Large Language Models (LLMs) by explicitly generating external chain-of-thought (COT) Wei et al. (2022) before producing final answers, achieving remarkable performance on complex reasoning tasks. Unlike traditional LLMs that primarily rely on intuitive, heuristic system 1 thinking, LRMs enhance their performance by employing analytical, deliberative system 2 thinking Qu et al. (2025); Li et al. (2025b), which includes sophisticated strategies such as self-reflection, multi-path exploration Li et al. (2025a); Yeo et al. (2025). Despite the benefits of system 2 thinking, a rigid reliance on this cautious thinking mode can also introduce new limitations, such as performance degradation in specific contexts Feng et al. (2025); Zhao et al. (2025) or overthinking Chen et al. (2024); Kumar et al. (2025a).
However, a research gap exists in thoroughly investigating how LRMs’ over-reliance on system 2 thinking affects their performance when confronting extremely simple questions better suited for intuition-driven system 1 processing. Specifically, existing research primarily analyzes LRM behaviors in complex tasks that naturally align with deliberative system 2 thinking, such as MATH for math problems Luo et al. (2025); Yang et al. (2025); Chen et al. (2024) or IFEval for instruction following Zhao et al. (2025). Although some benchmarks involve relatively simple questions for human Miao et al. (2021); Yan et al. (2025); Chiang and Lee (2024), these tasks are often adversarial or non-trivial for LLMs, requiring moderate reasoning and can not be effectively solved through intuitive system 1 processing. Thus, a benchmark to assess the system 1 thinking capability of LRMs is still lacking, further hindering our understanding of LRMs’ cognitive flexibility between the two systems Ziabari et al. (2025); Qu et al. (2025).
To fill this research gap, we introduce the System 1 Thinking Capability Benchmark (S1-Bench), which measures the performance of LRMs across various simple tasks that commonly encountered in real-world applications. S1-Bench has the following three characteristics: (1) Simple. The questions are not hard for humans and can be easily answered by LLMs. LLMs with 7-9B parameters can robustly provide correct answers through direct responses when sampled across multiple temperatures. (2) Diverse. S1-Bench is not limited to simple reasoning problems; it encompasses four major categories and 28 subcategories in two languages (English and Chinese), including commonsense knowledge, loosely constrained instruction following, and straightforward analytical problems. (3) Natural. The questions are clear, without any misleading elements or ambiguities, ensuring they can be answered intuitively.
S1-Bench conducted extensive evaluations across 22 LRMs, yielding the following key findings: (1) Current LRMs exhibit inefficiency and lack system 1 thinking capabilities, with average output lengths 15.5 times longer than small LLMs. (2) LRMs often arrive at correct answers early in their thinking process but continue unnecessary reasoning with gradually increasing content similarity. (3) Despite employing deep reasoning, LRMs exhibit accuracy degradation on simple questions compared to traditional LLMs. (4) LRMs show the ability to prejudge the simplicity of questions, yet they still exhibit inefficiency. These findings emphasize the significant distance LRMs must traverse to become powerful dual-system compatible models.
Our contributions can be summarized as follows:
  • To the best of our knowledge, S1-Bench is the first benchmark to evaluate the system 1 thinking capabilities of LRMs.
  • We introduce a workflow for constructing simple dataset suitable for system 1 evaluation.
  • Extensive experiments reveal the inefficiency and under-accuracy of LRMs on simple questions that are suited for system 1 thinking.
2 Related work
Figure 2: Construction workflow for S1-Bench and an illustrative example from each major category.
2.1 Large Reasoning Models
Large Reasoning Models (LRMs), characterized by explicitly generating external thinking processes before final answers Kumar et al. (2025b); Chen et al. (2025), achieve a paradigm shift from intuitive system 1 thinking to deliberative system 2 reasoning compared to traditional LLMs Li et al. (2025b); Qu et al. (2025), thus achieving superior performance on complex tasks. The development of recent LRMs has largely followed two main approaches: large-scale reinforcement learning (RL) and model distillation. Models trained via large-scale RL Guo et al. (2025); Team (2025b); Team et al. (2025) leverage reward-based optimization to gradually incentivize deliberative reasoning. In contrast, distillation-based LRMs OpenAI (2024); Min et al. (2024); Team (2025a); Ye et al. (2025); Muennighoff et al. (2025) acquire such abilities by transferring structured reasoning patterns from advanced teacher models.
2.2 Limitations of LRMs
While LRMs have shown significant performance gains through deliberate reasoning, rigid adherence to this overly cautious thinking can introduce new limitations. On one hand, intermediate reasoning steps can cause excessive token generation and unnecessary solving attempts Chen et al. (2024); Hashemi et al. (2025); Kumar et al. (2025a). On the other hand, LRMs’ performance can drop in specific contexts like safety scenarios Jiang et al. (2025); Zhao et al. (2025) and role-playing Feng et al. (2025). However, prior studies mainly evaluated LRMs on complex tasks that more suited for deliberative system 2 thinking. Our work examines how deliberative reasoning impacts extremely simple problems better matched to intuition-driven system 1 processing.
3 S1-Bench
We introduce S1-Bench, a bilingual, multi-domain benchmark designed to evaluate system 1 thinking capability of LRM on extremely simple questions. These questions are easily solvable by traditional LLMs and not hard for humans. S1-Bench, which covers both English and Chinese, is organized into four major categories: reasoning (Rsn), knowledge (Kno), instruction following (If) and analysis (Ana), representing major dimensions commonly employed in LLM capability evaluation Zheng et al. (2023); Chang et al. (2024).
This section begins with how simplicity is ensured, then the detailed construction workflow for S1-Bench, and concludes with an overview of the dataset statistics. Figure 2 shows the construction workflow and an illustrative example per category.
3.1 How to Ensure Simplicity?
We ensure questions are simple and suitable for system 1 thinking through two aspects.
3.1.1 A Priori Simplicity Constraints
We require all questions and answers to adhere to the following guidelines:
(1) Questions must be naturally, clearly expressed, and unambiguous, free of intentional traps.
(2) Answers must be unique or easily falsifiable (such as providing a three-letter English word).
We’ve further implemented category-specific constraints to ensure simplicity: Rsn: Limited to problems solvable with minimal reasoning or intuition. Kno: Restricted to common knowledge with unique, verifiable answers from sources like Wikipedia. If: Involve straightforward instructions without strict formatting requirements. Ana: Limited to questions whose answers can be directly inferred from the prompt, such as binary classification. These constraints ensure all questions remain straightforward for human respondents.
3.1.2 A Posteriori Simplicity Verification
Due to the biases existing between language models and humans Gallegos et al. (2024), questions that are simple for humans may be difficult for LLMs. Therefore, we introduce additional posteriori verification to ensure that questions are simple enough to be correctly and robustly answered by smaller LLMs from different families. This alleviates cognitive biases in traditional LLMs for system 1 thinking problems Li et al. (2025b).
3.2 Construction Workflow
To implement a priori constraints, we introduce data generators [^2] We select Claude-3.7-Sonnet and Qwen2.5-72B-Instruct. and quality discriminators [^3] We select GPT-4o and DeepSeek-V3-241226. that adhere to the a priori settings. To ensure posterior verification, we introduce simplicity validators [^4] We select four small LLMs: Qwen2.5-7B, Llama3.1-8B, Mistral-8B, and Gemma2-9B. The full model IDs are detailed in Table 4. and correctness evaluator [^5] We select GPT-4o.. The generators, discriminators, and evaluator are uniformly configured with temperature 0.7 and top-p=0.9, while validators are set with temperatures of 0.0, 0.2, and 0.4, each sampling 10 times. Additionally, we recruit three experienced graduate students as annotators, familiar with LLMs and fully comprehending the construction objectives of S1-Bench. For the prompts used in the workflow, refer to Table 11.
Subcategory Preparation. Initially, to ensure diversity, we draw on existing benchmark classification methods (e.g., MMLU, IFEval and GSM8K) and evaluation surveys Chang et al. (2024), selecting or merging existing subcategories, and designing new ones. We ensure that all subcategories meet the requirements for simplicity and provide an example question for each.
Generation and Modification. We use two data generators to create initial questions and answers for each candidate subcategory. The generation process adheres to prompt requirements (refer to Table 11), ensuring bilingual content that aligns with subcategory definitions while maintaining diverse perspectives. Each question is accompanied by auxiliary answers, with 50 initial bilingual question-answer pairs generated per subcategory. These pairs are independently assessed by three annotators and two quality discriminators , who ensure that the questions are unambiguous, have clear boundaries, and admit only one correct answer. For each question-answer pair, five independent assessments are conducted. Based on these evaluations, annotators discuss and collectively decide whether to retain, modify, or discard each question, ensuring alignment with the requirements outlined in Section 3.1.1. Retained questions are submitted to validators for testing, and an evaluator assesses the accuracy of responses. Questions with correct answers are included in S1-Bench, while others undergo reprocessing.
Questions with errors in any of the 30 sampling iterations undergo an iterative difficulty reduction process. These questions return to generators with specific prompts to decrease complexity, followed by the same discrimination, discussion, and evaluation procedures. Questions failing to achieve error-free performance after three difficulty reduction iterations are excluded from the workflow. The final S1-Bench comprises questions that satisfy both human-established a priori simplicity constraints and a posteriori simplicity verification.
Figure 3: Statistical distribution of token counts for S1-Bench questions.
Model t=0.0 t=0.2 t=0.4 Tokens
Gemma2-9B 100.00 100.00 100.00 38.77
Llama3.1-8B 100.00 100.00 100.00 42.00
Mistral-8B 100.00 100.00 100.00 44.38
Qwen2.5-7B 100.00 100.00 100.00 42.81
Qwen2.5-14B 99.74 99.76 99.76 40.00
Qwen2.5-32B 99.98 99.98 99.98 43.17
Qwen2.5-72B 100.00 100.00 100.00 44.61
Llama3.3-70B 100.00 99.76 99.76 53.71
DeepSeek-v3 100.00 100.00 100.00 79.53
Table 1: Average accuracy and response token count of different LLMs, each sampled 10 times at three temperature settings.
3.3 Benchmark Statistics
S1-Bench comprises 422 question-answer pairs across four major categories and 28 subcategories, balanced with 220 English and 202 Chinese questions. We present the category distribution and descriptions with examples for all subcategories, which can be found in Figure 8 and Table 15,16 in the appendix. Figure 3 shows the token length distribution, with questions averaging 14.46 tokens.
To validate the absence of response bias from validators, we tested S1-Bench on five LLMs that serve as base models for commonly used LRMs. As shown in Table 1, these models achieved an accuracy exceeding 99% under three temperature settings, confirming the simplicity and robustness of S1-Bench.
Model ID Size Loose Format Strict Format L-Corr $\uparrow$ S-Corr $\uparrow$ $\mathtt{ART}\downarrow$
pass@1$\uparrow$ acc@$\uparrow$ pass@1$\uparrow$ acc@$\uparrow$
Validator LLMs 7-9B 100.00 100.00 100.00 100.00 42.00
QwQ-32B 32B 100.00 100.00 100.00 100.00 100.00 100.00 720.10
Hunyuan-T1 99.91 99.53 99.91 99.53 100.00 100.00 542.31
Sky-T1-32B 32B 98.82 94.79 94.88 79.62 99.48 95.26 163.00
DS-R1 671B 100.00 100.00 100.00 100.00 100.00 100.00 646.40
DS-R1-70B 70B 99.48 97.39 99.38 96.92 100.00 99.91 453.81
DS-R1-32B 32B 99.72 98.82 99.72 98.82 100.00 100.00 429.91
DS-R1-14B 14B 99.57 97.87 99.57 97.87 100.00 100.00 475.46
DS-R1-8B 8B 97.44 97.16 97.39 97.16 99.76 99.53 452.11
DS-R1-7B 7B 95.21 85.78 95.21 85.78 99.24 99.24 454.55
DS-R1-1.5B 1.5B 81.47 54.50 81.47 54.50 97.58 97.58 489.54
Nemotron-49B 49B 99.15 97.39 99.15 97.39 100.00 100.00 362.54
Nemotron-8B 8B 86.16 69.91 79.81 59.00 99.43 84.31 372.57
L-R1-32B 32B 97.87 91.00 94.74 79.62 98.91 95.07 1095.36
L-R1-32B-DS 32B 99.57 98.10 99.57 98.10 99.81 99.81 524.12
L-R1-14B-DS 14B 99.05 95.97 99.05 95.97 99.19 99.19 693.19
L-R1-7B-DS 7B 94.64 83.65 94.64 83.65 99.76 99.67 496.47
s1.1-32B 32B 99.53 98.34 99.48 98.10 99.57 99.53 998.00
s1.1-14B 14B 97.63 93.60 97.25 91.94 97.77 97.39 839.86
s1.1-7B 7B 96.68 88.39 88.58 63.98 97.11 88.96 711.49
EXAONE-32B 32B 97.06 94.08 97.06 94.08 99.81 99.81 800.56
EXAONE-7.8B 7.8B 88.15 75.12 87.82 74.41 98.72 98.06 1046.87
EXAONE-2.4B 2.4B 72.42 56.16 72.32 56.16 97.44 97.25 1593.96
Table 2: Main results in the top-p sampling setting on the S1-Bench, sorted by model family. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst.
4 Main Experiment
4.1 Baseline Models and Configurations
We evaluated 22 different LRMs, which were explicitly trained to first respond with a thinking process, and then generate a final answer. These LRMs include open-source families, such as DeepSeek Guo et al. (2025), Qwen Team (2025b), Nemotron Chris Alexiuk and Patel (2025), Light-R1 Wen et al. (2025), s1.1 Muennighoff et al. (2025), EXAONE Research et al. (2025), and Sky-T1 Griggs et al. (2025), as well as closed-source Hunyuan-T1 Tencent (2025), spanning from tiny (1.5B) to large (671B) parameter sizes [^6] More details in Appendix Table 5 (includes training data scale and algorithms).. Notably, OpenAI’s o-series models are not included as they do not disclose thinking processes to users. For each model, we consider two sets of generation configurations: Greedy sampling with temperature t = 0; Top-p sampling with temperature t = 0.6, top-p = 0.95 and sampling size k=5.
4.2 Evaluation Metrics
Format Metrics.
Generally, the output of LRMs should be separated by an “end thinking marker” (e.g., </think>) that distinguishes the thinking process from the final answer. However, LRMs do not always generate responses in correct format. To assess this, we calculate the percentage of responses that meet the required formatting criteria, categorized as follows (for top-p sampling setting, each metric is averaged across 5 generations):
S-Corr (Strict Format Correctness Rate): The percentage of responses that strictly adhere to the format, meaning they only contain one "end thinking marker" followed by a non-empty final answer.
L-Corr (Loose Format Correctness Rate): The percentage of responses that loosely adhere to the format, meaning they meet the format requirements except for responses with endless repetitions [^7] Detailed format analyses are in the Appendix C..
Model ID Size S1-Bench-$\mathtt{EN}$ S1-Bench-$\mathtt{ZH}$ Avg
Rsn Kno If Ana Avg Rsn Kno If Ana Avg
Gemma2-9B 9B 74.8 29.4 5.3 52.4 45.9 51.6 19.8 7.5 35.1 31.0 38.8
Llama3.1-8B 8B 91.0 35.4 12.4 61.9 56.0 44.0 28.3 15.2 18.7 26.7 42.0
Qwen2.5-7B 7B 65.5 46.3 6.4 49.6 46.5 50.5 46.6 9.8 36.9 38.8 42.8
Mistral-8B 8B 67.2 55.5 8.6 50.1 49.6 47.3 56.1 14.8 29.7 38.7 44.4
Column Avg - 74.6 41.6 8.2 53.5 49.5 48.3 37.7 11.8 30.1 33.8 42.0
Sky-T1-32B 32B 215.8 174.1 98.5 233.3 194.3 125.5 125.3 99.4 145.5 128.9 163.0
Nemotron-49B 49B 599.7 587.6 396.5 526.1 540.4 232.9 157.3 235.5 107.8 168.8 362.5
Nemotron-8B 8B 561.0 585.1 458.0 303.1 462.6 369.5 326.0 288.1 166.7 273.5 372.6
DS-R1-32B 32B 421.8 504.4 414.7 521.1 473.7 362.2 385.6 343.1 408.8 382.2 429.9
DS-R1-8B 8B 472.2 528.9 530.7 462.7 491.2 521.9 404.4 266.2 395.5 409.4 452.1
DS-R1-70B 70B 464.1 501.3 378.5 536.1 484.0 450.8 450.2 328.4 416.7 420.9 453.8
DS-R1-7B 7B 447.5 623.9 353.8 510.0 495.5 446.5 463.2 339.5 373.0 409.4 454.5
DS-R1-14B 14B 503.7 674.7 367.3 494.2 519.0 452.0 465.4 375.3 405.8 428.0 475.5
DS-R1-1.5B 1.5B 480.8 584.7 417.4 577.2 529.1 493.0 497.4 329.8 423.1 446.0 489.5
L-R1-7B-DS 7B 568.1 667.1 501.7 566.3 580.3 444.8 454.6 344.1 366.4 405.0 496.5
L-R1-32B-DS 32B 574.5 706.6 647.6 632.8 636.3 431.2 367.0 377.1 418.7 402.2 524.1
Hunyuan-T1 561.6 693.8 380.9 435.0 521.2 676.8 553.8 505.1 523.8 565.3 542.3
DS-R1 671B 786.1 723.8 711.4 529.2 672.5 727.3 638.5 607.9 533.9 617.9 646.4
L-R1-14B-DS 14B 951.0 1026.0 829.8 653.5 848.2 594.7 610.1 442.2 451.7 525.7 693.2
s1.1-7B 7B 1039.5 840.8 1923.2 529.4 929.9 489.6 351.3 1034.3 332.4 475.6 711.5
QwQ-32B 32B 873.3 808.1 520.8 634.7 722.4 866.9 707.3 613.3 667.7 717.6 720.1
EXAONE-32B 32B 1323.7 1057.6 1537.0 711.6 1086.4 703.2 348.6 1302.9 125.5 490.3 800.6
s1.1-14B 14B 871.8 746.2 2233.1 708.1 960.2 654.6 546.0 1512.6 579.7 710.7 839.9
s1.1-32B 32B 1077.9 889.7 2055.4 781.7 1081.7 995.6 765.2 1634.6 666.5 906.5 998.0
EXAONE-7.8B 7.8B 1498.3 1398.9 1775.7 882.4 1303.8 1410.3 497.8 1633.1 205.0 767.0 1046.9
L-R1-32B 32B 1614.0 1217.3 1996.9 930.1 1338.3 1035.6 737.7 1240.7 610.2 835.3 1095.4
EXAONE-2.4B 2.4B 1927.3 1426.2 1200.1 825.7 1320.7 2469.7 1622.6 2471.6 1511.2 1898.7 1594.0
Column Avg - 810.6 771.2 896.8 590.2 736.0 679.8 521.6 742.0 447.1 563.0 652.8
Improvement - $\times$10.9 $\times$18.5 $\times$109.7 $\times$11.0 $\times$14.9 $\times$14.1 $\times$13.8 $\times$62.7 $\times$14.8 $\times$16.7 $\times$15.5
Table 3: $\mathtt{ART}$ on the S1-bench across two languages and four main categories. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst. Bold represents the maximum Improvement value for each language.
Efficiency Metrics.
We calculate the Average Response tokens (ART), which represent the average number of tokens in responses that meet the loose formatting requirements. Token counts are obtained using the Qwen2.5 tokenizer.
Accuracy Metrics.
We calculate accuracy metrics for both strict and loose formatting requirements, respectively. We use GPT-4o as the evaluator to assess the correctness of the responses [^8] If a final answer can be isolated, only the final answer is evaluated; otherwise, the entire response is assessed., with the evaluation prompt in Table 11. For greedy sampling, we directly calculate the accuracy rate, with complete results in Appendix D. For top-p sampling, we utilize two metrics:
Pass@1: Followed DeepSeek-R1 Guo et al. (2025), we calculate pass@1 to assess the percentage of correct responses among the k=5 generations. Specifically, it is defined as:
[…]
where $p_{i}$ is the correctness of the i-th generation.
Acc@k: Since S1-Bench is composed of extremely simple questions, we calculate acc@k. Specifically, acc@k=1 if all k responses are correct and acc@k = 0 otherwise. It is defined as:
[…]
Notably, S-Corr represents the upper bound for pass@1 and acc@5 under strict formatting requirements, with L-Corr following the same principle.
4.3 Main Results
Table 2 and Figure 1 presents the main results of LRMs on S1-Bench, revealing two key phenomena.
LRMs exhibit significantly lower efficiency than LLMs on S1-Bench, and no clear correlation is observed between ART and model size.
Notably, state-of-the-art LRMs, such as DeepSeek-R1 and QwQ-32B, do not demonstrate a distinct advantage in efficiency. In contrast, Sky-T1-32B, which undergoes specific optimizations to mitigate overthinking using SimPO, achieves the highest efficiency. The L-R1-DS 7B/14B/32B models generate longer responses than the DS-R1-7B/14B/32B models, and the former are further post-trained on the latter. This suggests that further post-training may improve reasoning ability for complex tasks, but at the cost of response efficiency for simple tasks. Finally, the s1.1 models generate considerably longer responses than the DeepSeek-R1-Distilled models. Despite both models being trained solely with SFT to acquire long-COT reasoning ability, the DeepSeek-R1-Distilled models use 800K training samples, while the s1.1 models are trained on only 1K. This discrepancy suggests that the smaller training set may lead to superficial imitation of long reasoning patterns, resulting in verbose thinking on simple questions.
Several LRMs exhibit under-accuracy on simple questions.
Our observations reveal that, despite employing deep reasoning, most LRMs tend to exhibit lower accuracy on simple questions compared to traditional LLMs. For example, DS-R1-1.5B and EXAONE-2.4B achieve just above 50% acc@k. Moreover, accuracy decreases with smaller model size. Finally, many LRMs struggle with robust correctness in top-p sampling, where acc@k is significantly lower than pass@1. This issue is particularly pronounced in smaller LRMs. For instance, DS-R1-1.5B achieved 81.47% pass@1 but only 54.50% acc@k.
5 Efficiency Analysis
Figure 4: (a) Comparison of Initial and Additional Thinking Costs for Each LRM. (b) Distribution of Solution Rounds for Each LRM.
5.1 ART Analysis of Question Types
To better understand the efficiency differences of LRMs across question types, we analyze $\mathtt{ART}$ across 4 main categories, 28 subcategories, and two languages, yielding the following insights.
LRMs exhibit the most significant increase in ART for instruction-following questions and tend to over-explore when the solution space is vast.
Although small LLMs generate the most concise responses for instruction-following questions across all main categories in S1-Bench, LRMs produce responses that are 109.7× longer in Chinese and 62.7× longer in English, making them the longest response type across all categories. We further analyze the subcategories of instruction-following questions. As shown in Figure 9, $\mathtt{ART}$ is notably longer in the subcategories of length constraints, character constraints, and sentence constraints. These three question types share a similar characteristic: their correctness is verifiable, but the solution space is vast. We find that, although the model quickly identifies a correct answer, it becomes trapped in the search space, continually exploring alternatives and failing to stop in time. This phenomenon is more pronounced in families with higher $\mathtt{ART}$, such as s1.1 and the EXAONE.
LRMs also demonstrate a significant increase in ART for reasoning tasks.
In Table 3, $\mathtt{ART}$ for reasoning questions also increases significantly, and Figure 9 further corroborates this phenomenon in $\mathtt{ART}$ in various reasoning subcategories. One possible explanation for this is that reasoning questions align closely with the post-training data distribution of LRMs, thereby further stimulating the long-chain reasoning behavior of LRMs.
LRMs consistently exhibit inefficiency in both Chinese and English.
In table 3, most LRMs show higher $\mathtt{ART}$ on English questions, while QwQ-32B and Hunyuan-T1 exhibit comparable $\mathtt{ART}$ across Chinese and English.
Figure 5: Maximum similarity between each segment and all preceding segments for LRMs across four categories.
5.2 Solution Analysis in Thinking Processes
To better understand the causes of inefficiency in LRMs on the S1-Bench, we analyze the thinking process where the final answer is correct and the format is strictly correct and non-empty [^9] We selected this subset because incorrect answers make it unclear whether LRMs are over-reasoning or under-reasoning, and malformed or empty thinking processes cannot be precisely extracted.. We begin by segmenting each thinking process into several solutions, each defined as a point at which LRMs explicitly arrive at a conclusion that directly aligns with the correct answer. The segmentation process is performed by DeepSeek-v3, with prompts detailed in Table 12. We then compute the average initial thinking cost for LRMs. For each sample, if the thinking process contains at least one solution, the cost is defined as the tokens counts in the first solution. If no clear and correct solution is provided, the cost is the total token counts in the thinking process. The reported value is the average across all selected samples. We find that:
The step-by-step reasoning in the initial thinking process is not the primary cause of the inefficiency in LRMs.
As shown in Figure 4 (a), although the total counts of thinking tokens varies considerably across LRMs, their initial thinking costs are similar and account for only a small fraction of the total.
Generating unnecessary solution rounds after reaching the correct answer is one of the reasons for the inefficiency of LRMs.
We further examine the distribution of solution rounds among various LRMs on the S1-Bench (Figure 4 (b)) and find that models with longer thinking processes tend to produce excessive solution rounds, repeatedly re-verifying simple problems that have already been solved. This redundant verification significantly contributes to computational inefficiency.
Figure 6: Distribution of the thinking process across four categories. FA and TP refer to Final Answer and Thinking Process, respectively. Green bars indicate cases where the final answer is correct, while red bars indicate cases where it is incorrect.
5.3 Redundancy in Thinking Processes
Information redundancy increases as reasoning sequences increase. We conduct a similarity analysis to analyze how information redundancy in the thinking processes of LRMs changes as reasoning sequences increase. Specifically, we first divide the complete thinking process into k equal-length segments [^10] We set k=15, changing its value does not affect the conclusions.. Then, we encode each segment using the all-MiniLM-L6-v2 model [^11] https://huggingface.co/sentence-transformers/allMiniLM-L6-v2. For each segment, we calculate the cosine similarity with all its preceding segments and use the maximum similarity as a measure of its information redundancy. As shown in Figure 5, information redundancy increases across all four main categories as reasoning sequences increase. Sky-T1-32B shows overall lower similarity, which stems from its shorter thinking process, but still demonstrates an upward trend.
6 Analysis of Errors
Figure 7 explores the relationship between thinking process quality and final answer accuracy. We focus on samples whose format is strictly correct and thinking process is non-empty. For samples with correct final answers, we categorize them based on whether the thinking process contains explicit incorrect conclusions in intermediate steps. For samples with incorrect final answers, we categorize them based on whether the correct answer is mentioned at least once during reasoning. We use DeepSeek-v3 for categorization, with the prompt provided in Table 13. Figure 6 presents the distribution of these four categories in different RLMs. Our analysis yields the following observations: (1) LRMs with lower accuracy often include incorrect intermediate conclusions in their reasoning, even when they ultimately reach correct final answers (light green). (2) Although LRMs sometimes reach the correct answer during reasoning, they may deviate and ultimately produce incorrect conclusions (light red).
7 Simplicity Prejudgement
Figure 7: Counting instances of prejudgment generated by LRMs for simple questions and $\mathtt{ART}$.
Finally, we discover an intriguing phenomenon: LRMs can prejudge certain simple questions. Specifically, we extract the first 50 tokens from each thinking process and use DeepSeek-v3 to identify sentences showing judgment, further extracting judgment words [^12] An example of prejudgment: “This question seems simple”. For prompt references, see Table 14.. Figure 7 displays the frequency of words judging questions as simple, with details provided in the Table 7. Our findings reveal:
LRMs possess the ability to prejudge question simplicity, especially in Chinese.
All LRMs exhibit prejudgment phenomena within their thinking processes, demonstrating an ability to directly assess question difficulty. Additionally, 14 LRMs exhibit significantly stronger prejudgment tendencies in Chinese, with this phenomenon concentrating primarily within the L-R1, DS-R1, QwQ, and Hunyuan model families.
Even with prejudgment, the thinking length of LRMs does not shorten.
As shown in Figure 7, the average $\mathtt{ART}$ of the thinking process exhibiting prejudgment does not decrease. We consider further exploration of this phenomenon as an important direction for future work.
These results suggest LRMs possess an inherent understanding of question difficulty. This opens a novel pathway toward dual-system compatibility for LRMs, which we identify as future work.
8 Conclusion
This paper presents S1-Bench, designed to evaluate LRMs’ performance on simple questions suited for system 1 thinking. Experiments reveal significant overthinking and performance degradation by LRMs on S1-Bench, highlighting the gap between current LRMs and dual-system compatibility.
Appendix A Subcategories in S1-Bench
Figure 8: S1-Bench Category Display. The inner circle represents four major categories, and the outer circle includes 28 subcategories.
Figure 8 shows the pie chart distribution of 28 subcategories in S1-Bench. For more details on the subcategories, please refer to Table 15,16.
Appendix B Baseline Model Details
Model Model ID URL
Qwen2.5-7B Qwen2.5-7B-Instruct https://huggingface.co/Qwen/Qwen2.5-7B-Instruct
Llama3.1-8B Llama-3.1-8B-Instruct https://huggingface.co/meta-llama/Llama-3.1-8B-Instruct
Mistral-8B Ministral-8B-Instruct-2410 https://huggingface.co/mistralai/Ministral-8B-Instruct-2410
Gemma2-9B gemma-2-9b-it https://huggingface.co/google/gemma-2-9b-it
Qwen2.5-14B Qwen2.5-14B-Instruct https://huggingface.co/Qwen/Qwen2.5-14B-Instruct
Qwen2.5-32B Qwen2.5-32B-Instruct https://huggingface.co/Qwen/Qwen2.5-32B-Instruct
Qwen2.5-72B Qwen2.5-72B-Instruct https://huggingface.co/Qwen/Qwen2.5-72B-Instruct
Llama3.3-70B Llama-3.3-70B-Instruct https://huggingface.co/meta-llama/Llama-3.3-70B-Instruct
DeepSeek-v3 DeepSeek-V3-0324 https://huggingface.co/deepseek-ai/DeepSeek-V3-0324
Table 4: Mapping of LLM abbreviations and IDs used in this paper, with their open-source URLs.
Model ID Abbreviation Base Model Alg. Size
DeepSeek
DeepSeek-R1-Distill-Qwen-1.5B DS-R1-1.5B Qwen2.5-Math-1.5B SFT 800K
DeepSeek-R1-Distill-Qwen-7B DS-R1-7B Qwen2.5-Math-7B SFT 800K
DeepSeek-R1-Distill-Llama-8B DS-R1-8B Llama-3.1-8B SFT 800K
DeepSeek-R1-Distill-Qwen-14B DS-R1-14B Qwen2.5-14B SFT 800K
DeepSeek-R1-Distill-Qwen-32B DS-R1-32B Qwen2.5-32B SFT 800K
DeepSeek-R1-Distill-Llama-70B DS-R1-70B Llama-3.3-70B-Instruct SFT 800K
DeepSeek-R1 DS-R1 DeepSeek-V3-0324 SFT&RL 800K&–
Qwen
QwQ-32B QwQ-32B Qwen2.5-32B
qihoo360
Light-R1-7B-DS L-R1-7B-DS DeepSeek-R1-Distill-Qwen-7B SFT 3K
Light-R1-14B-DS L-R1-14B-DS DeepSeek-R1-Distill-Qwen-14B SFT&RL 3K&–
Light-R1-32B-DS L-R1-32B-DS DeepSeek-R1-Distill-Qwen-32B SFT 3K
Light-R1-32B L-R1-32B Qwen2.5-32B-Instruct SFT&DPO 73K&–
simplescaling
s1.1-7B s1.1-7B Qwen2.5-7B-Instruct SFT 1K
s1.1-14B s1.1-14B Qwen2.5-14B-Instruct SFT 1K
s1.1-32B s1.1-32B Qwen2.5-32B-Instruct SFT 1K
LG AI Research
EXAONE-Deep-2.4B EXAONE-2.4B EXAONE-3.5-2.4B-Instruct SFT&DPO&RL 1.6M&20K&10K
EXAONE-Deep-7.8B EXAONE-7.8B EXAONE-3.5-7.8B-Instruct SFT&DPO&RL 1.6M&20K&10K
EXAONE-Deep-32B EXAONE-32B EXAONE-3.5-32B-Instruct SFT&DPO&RL 1.6M&20K&10K
NVIDIA
Llama-3.1-Nemotron-Nano-8B-v1 Nemotron-8B Llama-3.1-8B-Instruct SFT&RL –&–
Llama-3.3-Nemotron-Super-49B-v1 Nemotron-49B Llama-3.3-70B-Instruct SFT&RL –&–
NovaSky
Sky-T1-32B-Flash Sky-T1-32B Qwen2.5-32B-Instruct SFT&SimPO 17K&10K
Table 5: The open-source LRMs details evaluated for S1-Bench.
Table 4 presents the abbreviations, IDs, and URLs of LLMs used in this paper. Table 5 displays the abbreviations, IDs, URLs, organizations, algorithms, and training data volumes of open-source LRMs evaluated in this study.
Appendix C Format Details
Format ID marker marker marker thinking final
(standard) (special) (number) process answer
Strict 100 $\surd$ 1 $\surd$ $\surd$
101 $\surd$ 1 $\times$ $\surd$
Loose (w/o Strict) 200 $\surd$ 1 $\surd$ $\times$
201 $\surd$ >1 $\surd$ $\surd$
202 $\surd$ >1 $\surd$ $\times$
203 $\surd$ >1 $\times$ $\surd$
204 $\times$ $\surd$ $\geq$1 $\surd$ $\surd$
205 $\times$ $\surd$ $\geq$1 $\surd$ $\times$
206 $\times$ $\surd$ $\geq$1 $\times$ $\surd$
207 $\times$ $\times$ 0 $\surd$
Error 300 $\surd$ $\surd$ $\geq$1 $\times$ $\times$
301 $\times$ $\times$ 0 $\infty$
Table 6: Format types.
Words Count
{CJK}UTF8gkai简单 466
straightforward 231
simple 118
{CJK}UTF8gkai不难 72
{CJK}UTF8gkai基础的 44
basic 24
{CJK}UTF8gkai直接 15
{CJK}UTF8gkai常见 10
{CJK}UTF8gkai基础 10
{CJK}UTF8gkai不复杂 7
simplest 7
pretty straightforward 5
easy 5
{CJK}UTF8gkai基本的 5
{CJK}UTF8gkai基本 3
Table 7: Top 15 phrases with prejudgment phenomenon, with counts summed across 22 LRMs.
Model Loose Format Error Format
Strict Format
100 101 200 201 202 203 204 205 206 207 300 301
QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
hunyuan-t1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Sky-T1-32B 95.26 0.00 0.62 0.09 0.19 0.00 0.28 0.00 0.00 3.03 0.00 0.52
DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-70B 99.91 0.00 0.09 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-14B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
DS-R1-7B 99.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.76
DS-R1-1.5B 97.58 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.42
Nemotron-49B 66.07 33.93 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Nemotron-8B 58.06 26.26 0.00 0.00 0.00 0.09 0.00 0.00 0.00 15.02 0.00 0.57
L-R1-32B 95.07 0.00 0.00 0.00 0.00 0.00 0.81 0.00 0.00 3.03 0.00 1.09
L-R1-32B-DS 99.81 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
L-R1-14B-DS 99.19 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.81
L-R1-7B-DS 99.67 0.00 0.05 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
s1.1-32B 99.53 0.00 0.00 0.05 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.43
s1.1-14B 97.39 0.00 0.00 0.14 0.00 0.00 0.24 0.00 0.00 0.00 0.00 2.23
s1.1-7B 88.96 0.00 0.00 7.96 0.09 0.00 0.00 0.00 0.00 0.09 0.00 2.89
EXAONE-32B 67.39 32.42 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.19
EXAONE-7.8B 65.83 32.23 0.00 0.00 0.05 0.47 0.00 0.00 0.00 0.14 0.00 1.28
EXAONE-2.4B 81.42 15.83 0.00 0.09 0.00 0.05 0.00 0.00 0.00 0.05 0.00 2.56
Table 8: Format types rates in top-p sampling setting.
Model Loose Format Error Format
Strict Format
100 101 200 201 202 203 204 205 206 207 300 301
QwQ-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
hunyuan-t1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Sky-T1-32B 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.47 0.00 0.24
DS-R1 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-70B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-32B 100.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
DS-R1-14B 99.76 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24
DS-R1-8B 99.53 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.24 0.00 0.24
DS-R1-7B 97.87 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 2.13
DS-R1-1.5B 91.94 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 8.06
Nemotron-49B 60.90 39.10 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
Nemotron-8B 55.21 26.78 0.00 0.00 0.00 0.00 0.00 0.00 0.00 16.35 0.00 1.66
L-R1-32B 85.55 0.24 0.00 0.24 0.71 0.24 0.95 0.00 0.00 6.64 2.61 2.84
L-R1-32B-DS 99.29 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.71
L-R1-14B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
L-R1-7B-DS 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
s1.1-32B 98.82 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 1.18
s1.1-14B 95.97 0.00 0.00 0.24 0.00 0.00 0.24 0.00 0.00 0.00 0.00 3.55
s1.1-7B 87.91 0.00 0.00 6.64 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.45
EXAONE-32B 65.88 33.89 0.00 0.24 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
EXAONE-7.8B 63.51 33.65 0.00 0.00 0.00 0.24 0.00 0.00 0.00 0.24 0.00 2.37
EXAONE-2.4B 78.91 15.88 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 5.21
Table 9: Format types rates in greedy decoding setting.
C.1 Division of Different Format
This section presents a comprehensive taxonomy of format errors and emphasizes the importance of addressing these issues in future research endeavors. Unlike conventional LLMs, LRMs frequently exhibit format errors, manifested when responses fail to adhere to the requirement of generating a single End Thinking Marker ($\mathtt{ETM}$). This inconsistency creates significant challenges in distinguishing between the thinking process and the final answer. Format errors represent a distinct category from over-thinking and under-accuracy phenomena, yet similarly indicate vulnerability to elementary reasoning traps.
As illustrated in Table 6, we have identified and classified 12 distinct format types, each assigned a unique ID. These format types can be broadly categorized into Strict Format and Loose Format, with Strict Format representing a specialized subset of Loose Format.
Loose Format. ID-100 and ID-101 represent two standard decoding modes, categorized as Strict Format. Both contain only one $\mathtt{ETM}$. Among these, ID-100 includes a thinking process, while ID-101 only generates the $\mathtt{ETM}$ and final answer.
[…]
Error Format. LRMs produce two types of undecodable errors: ID-300 indicates that the LRM outputs only a $\mathtt{ETM}$ without any additional content, while ID-301 represents cases where the LRM’s thinking process reaches the maximum length.
C.2 Evaluation of Different Format
In the classification shown in Table 6, ID-100 / 101 / 201 / 203 / 204 / 206 / 207 have decomposable final answers. If multiple $\mathtt{ETM}$s appear, only the last one is selected for splitting. In this case, the evaluator uses the final answer as the evaluation target.
ID-200 / 202 / 205 lack a final answer, in which case the evaluation target becomes the thinking process, as this part may also contain the correct solution to the question.
C.3 Statistics of Different Format
Table 8 presents the distribution of 12 format types under top-p sampling settings. Our analysis reveals three key insights: (1) The infinite thinking phenomenon is widespread across most models, particularly concentrated in LRMs with fewer than 32B parameters. This indicates that output formatting should be prioritized during training to reduce format errors, especially when working with less powerful foundation models. (2) The Nemotron and EXAONE families frequently produce correctly formatted responses without visible reasoning processes. This behavior can be viewed as a mechanism for mitigating over-thinking. However, the EXAONE family still exhibits substantial overthinking tendencies, suggesting that LRMs’ capability to respond without visible reasoning and their tendency to overthink may be orthogonal characteristics. (3) None of the evaluated LRMs exhibited behaviors classified as ID-205/206, possibly because special $\mathtt{ETM}$s are perceived by LRMs as the only alternative to standard $\mathtt{ETM}$s. Furthermore, ID-300 was not observed, indicating that high-temperature sampling generally yields semantically meaningful responses.
Table 9 shows the format statistics of greedy sampling. The phenomenon is generally similar to top-p sampling. L-R1-32B produced ID-300, possibly due to poor compatibility between the training method and low temperature.
Appendix D Experiment Details
Table 10 shows LRMs’ performance on S1-Bench under greedy sampling settings. While overall accuracy improves compared to top-p sampling settings, $\mathtt{ART}$ slightly decreases, yet overthinking and accuracy degradation issues persist. Figure 9 illustrates the $\mathtt{ART}$ heatmap across 28 subcategories and 26 models.
We compile the statistics of the prejudgment words extracted, as shown in Table 7.
Model Size acc (Loose) acc (Strict) $\mathtt{SUC}$-$\mathtt{L}\uparrow$ $\mathtt{SUC}$-$\mathtt{S}\uparrow$ $\mathtt{ART}$downarrow
QwQ-32B 32B 100.00 100.00 100.00 100.00 750.41
Hunyuan-T1 100.00 100.00 100.00 100.00 541.09
Sky-T1-32B 32B 99.53 99.05 99.76 99.29 157.12
DS-R1 671B 100.00 100.00 100.00 100.00 621.89
DS-R1-70B 70B 99.76 99.76 100.00 100.00 469.78
DS-R1-32B 32B 100.00 100.00 100.00 100.00 428.46
DS-R1-14B 14B 99.29 99.29 99.76 99.76 463.52
DS-R1-8B 8B 97.63 97.39 99.76 99.53 452.11
DS-R1-7B 7B 94.31 94.31 97.87 97.87 436.87
DS-R1-1.5B 1.5B 76.54 76.54 91.94 91.94 473.67
Nemotron-49B 49B 99.53 99.53 100.00 100.00 337.94
Nemotron-8B 8B 84.60 77.73 98.34 81.99 446.62
L-R1-32B 32B 92.18 85.78 94.55 85.78 996.36
L-R1-32B-DS 32B 99.29 99.29 99.29 99.29 528.45
L-R1-14B-DS 14B 98.82 98.82 98.82 98.82 664.28
L-R1-7B-DS 7B 92.65 92.65 98.82 98.82 514.60
s1.1-32B 32B 98.82 98.82 98.82 98.82 983.38
s1.1-14B 14B 95.97 95.50 96.45 95.97 786.30
s1.1-7B 7B 94.31 87.68 94.55 87.91 630.52
EXAONE-32B 32B 97.63 97.39 100.00 99.76 746.89
EXAONE-7.8B 7.8B 86.73 86.49 97.63 97.16 947.92
EXAONE-2.4B 2.4B 72.99 72.99 94.79 94.79 1394.72
Table 10: Main results in the greedy decoding setting on the S1-Bench, sorted by model family. Bold teal marks best performance, teal second best, bold burgundy worst, and burgundy second worst.
Figure 9: $\mathtt{ART}$ on the 28 subcategories, which is the average result of five generations under top-p sampling with Loose Format setting.
Appendix E Prompts discussed in this paper
This section presents the prompts used in this paper. For S1-Bench construction-related prompts, please refer to Table 11; for experimental prompts, please refer to Table 12,13 and 14.
Prompt for construction workflow for S1-Bench
Initial Generation
Based on the category name and explanation, generate 50 pairs of questions and answers in both Chinese and English. Ensure that the generated questions vary in perspective and content, maintaining diversity. All questions must have a **uniquely determined, unambiguous answer with no potential conditions causing multiple possible answers**. Ensure that the questions are expressed clearly. Additionally, the difficulty level of the questions should be **as simple as possible—so easy that they can be answered correctly without any need for thought**.
# Category Name:
{category}
# Explanation:
{explanation}
# Cases:
## English question:
{question_en}
## English Answer:
{answer_en}
## Chinese question:
{question_zh}
## Chinese Answer:
{answer_zh}
Please generate 50 pairs of Chinese and English questions and answers in the following format: [question]English-question[answer]English-answer[question]Chinese-question[answer]Chinese-answer…
Start generating:
Discriminating Generation Quality
Given a question, its answer, and its category, please analyze from the following perspectives as comprehensively as possible:
1. Whether the question belongs to the specified category.
2. Whether the question is clear, unambiguous, has an absolutely unique answer, and has no potential alternative interpretations.
3. Whether the answer is absolutely correct; if not, what the correct answer should be.
4. Whether the question is similar to other given questions, and if similar, whether more diverse questions can be generated.
# Question Category and Explanation:
{category_with_exp}
# Question and Answer:
{question_with_answer}
# Other Questions:
{questions_list}
Begin your analysis, aiming to be as detailed and comprehensive as possible:
Reduce Difficulty Generation
Given a question and answer that are too complex for the model to answer correctly, you need to further reduce their difficulty while trying to:
- Maintain the original perspective and semantic information
- Preserve the question’s alignment with its category
# Question Category and Explanation:
{category_with_exp}
# Question and Answer:
{question_with_answer}
The new question and answer:
Evaluation on S1-Bench
**Question:**
{question}
**Model Answer:**
{model_answer}
**Ground Truth:**
{ground_truth}
Your task is to evaluate whether the model’s answer is correct.
An answer is considered correct as long as it contains the ground truth (regardless of how complex or detailed the description is).
If there are parenthetical notes after the ground truth, then there may be multiple correct answers. In this case, the given answer is just one example, and any answer that meets the requirements specified in the notes can be considered correct.
Additionally, some reasonably uncertain supplementary information is also considered appropriate, including more details, possibilities, and expanded discussion. You should focus more on whether the reply contains the correct answer.
You need to output a standard JSON, providing your explanation of the evaluation in the "explain" field, and giving the evaluation result in the "result" field, where 1 means the answer is correct and 0 means it is incorrect.
Your action should follow the given format: "explain": "", "result": 0/1
Table 11: Prompt for construction workflow for S1-Bench. The “explanation” (and “category_with_exp”) includes the definition of each major category, simplicity requirements in Section 3.1.1, and the definition of subcategories.
Prompts for Solution Segmentation
# Task Description:
Your task is to segment the given Chain of Thought according to the following rules:
1. Segmentation positions:
(1) Please identify and extract all sub-reasoning processes from the Chain of Thought that meet the following condition: They explicitly arrive at a conclusion (including cases phrased as questions, e.g., "right?") that is directly consistent with the Ground Truth. Reasoning processes that only indirectly support the Ground Truth or result in partially aligned conclusions should be excluded.
(2) After clearly reaching the conclusion consistent with the Ground Truth, insert the segmentation marker <split> to distinctly separate each qualifying sub-reasoning process.
2. Output Restriction:
(1) You should only directly output the segmentation result without adding any additional supplements.
(2)Except for inserting the <split> separator, you must not make any other modifications to the original Chain of Thought, not even minor character-level changes such as punctuation, spacing, or capitalization. In other words, after removing all <split> separators, the output text must match the original Chain of Thought exactly.
# Some cases:
## Case 1:
### Question:
What is the value of x if x plus 3 equals 8?
### Ground Truth:
5
### Chain of Thought:
Okay, so I need to figure out what x is if x plus 3 equals 8. The problem says "x plus 3 equals 8," so in mathematical terms, that would be: x + 3 = 8
Alright, now I need to solve for x. From what I remember in algebra, the goal is to isolate x on one side of the equation. That means I need to get rid of that "+ 3" that’s being added to x. To do that, I should perform the opposite operation. Since it’s adding 3, the opposite would be subtracting 3.
So, if I subtract 3 from both sides, that should keep things equal. Let me try that: x + 3 - 3 = 8 - 3
Simplifying both sides, on the left side, the +3 and -3 cancel each other out, leaving just x. And on the right side, 8 minus 3 is 5. So that simplifies to: x = 5
Wait, let me check if that makes sense. If I substitute x back into the original equation, does it hold true? Let’s see:
Original equation: x + 3 = 8
Substitute x = 5: 5 + 3 = 8
5 plus 3 is indeed 8, so that works out.
Another way might be to consider what number, when you add 3 to it, gives 8. So starting from 8, if I take away 3, I should get back to the original number. That’s the same as subtracting 3 from 8, which again gives 5. So that’s consistent with the first method.
Wait a second, what if the original problem was written differently? The problem says "x plus 3 equals 8." In symbols, that’s unambiguous: x + 3 = 8. So no, I think that’s right.
*Final Answer**
The value of x is 555.
### Output:
Okay, so I need to figure out what x is if x plus 3 equals 8. The problem says "x plus 3 equals 8," so in mathematical terms, that would be:x + 3 = 8
Alright, now I need to solve for x. From what I remember in algebra, the goal is to isolate x on one side of the equation. That means I need to get rid of that "+ 3" that’s being added to x. To do that, I should perform the opposite operation. Since it’s adding 3, the opposite would be subtracting 3.
So, if I subtract 3 from both sides, that should keep things equal. Let me try that: x + 3 - 3 = 8 - 3
Simplifying both sides, on the left side, the +3 and -3 cancel each other out, leaving just x. And on the right side, 8 minus 3 is 5. So that simplifies to: x = 5<split>
Wait, let me check if that makes sense. If I substitute x back into the original equation, does it hold true? Let’s see:
Original equation: x + 3 = 8
Substitute x = 5: 5 + 3 = 8
5 plus 3 is indeed 8, so that works out.<split>
Another way might be to consider what number, when you add 3 to it, gives 8. So starting from 8, if I take away 3, I should get back to the original number. That’s the same as subtracting 3 from 8, which again gives 5.<split> So that’s consistent with the first method.
Wait a second, what if the original problem was written differently? The problem says "x plus 3 equals 8." In symbols, that’s unambiguous: x + 3 = 8. So no, I think that’s right.
*Final Answer**
The value of x is 555.<split>
## Case 2:
## Case 3:
# Eval Target:
## Question:
{question}
## Ground Truth:
{ground_truth}
## Chain of Thought:
{thinking_process}
## Output:
Table 12: Prompts for Solution Segmentation.
Prompts for Error Analysis
Types of Correct
# Task Description:
You will receive a Question, its corresponding Ground Truth, and a Chain of Thought(COT) generated by a LLM for that Question.
Your task is to carefully analyze the CoT and assign it to one of the two predefined categories listed below.
# Categories:
1: The CoT ***includes explicit incorrect conclusions*** in intermediate reasoning steps.
2: The CoT ***doesn’t include any explicit incorrect conclusion*** in intermediate reasoning steps.
Output your evaluation in the following format:
# TheReason: [note: Conduct a step-by-step analysis, stating if and where explicit incorrect conclusions occur in the COT.]
# ErrorType: [note: Summarize each incorrect conclusion into a specific error type using a phrase of less than 5 words, such as factual inaccuracies, logical fallacies, comprehension mistakes, calculation errors, formatting issues, and so forth, to better conduct further evaluation and analysis. Directly output a Python list, where each element represents the error type of a specific incorrect conclusion in the CoT. If there are no incorrect conclusions, return an empty list.]
# TheCategory: [note: Provide your classification based on your analysis using only the number "1" or "2". Do not add any additional text.]
# Question:
{question}
# Ground Truth:
{gound_truth}
# COT:
{thinking_process}
# TheReason:
Types of Error
# Task Description:
You will receive a Question, its corresponding Ground Truth, and a Chain of Thought(COT) generated by a LLM for that Question.
Your task is to carefully analyze the CoT and assign it to one of the two predefined categories listed below.
# Categories:
1: Regardless of whether the CoT ultimately arrives at the correct final answer or not, ***the correct answer is explicitly mentioned at least once*** within the reasoning steps (even if it is not ultimately adopted).
2: ***The correct answer is never explicitly mentioned or referenced*** at any point within the reasoning steps.
Output your evaluation in the following format:
# TheReason: [note: Conduct a step-by-step analysis, explicitly stating whether and where a correct answer is mentioned within the reasoning steps.]
# TheCategory: [note: Provide your classification based on your analysis using only the number "1" or "2". Do not add any additional text.]
# Question:
{question}
# Ground Truth:
{answer}
# COT:
{thinking_part}
# TheReason:
Table 13: Prompts for Error Analysis.
Prompts for Extracting Prejudgment Sentences and Words
Extracting Sentences
# Given the beginning of a model’s thought process when answering a question, determine if there is any evaluation by the model regarding the difficulty level or category of the question.
# For example:
Example Evaluation 1: This seems to be an easy question.
Example Evaluation 2: {CJK}UTF8gkai这看起来似乎是个基础的问题
Example Evaluation 3: This question is a bit complex.
Example Evaluation 4: {CJK}UTF8gkai这个问题看起来复杂
# Beginning of thought process:
{thinking_process}
# Please output only the evaluative short phrase you find. If there is no evaluation, output "None".
# Output:
Extracting Words
# You are given a comment sentence, and your goal is to find the only comment-related word.
# For example:
## case 1:
Sentence: {CJK}UTF8gkai这是一个很简单的算术题
Word: {CJK}UTF8gkai简单
## case 2:
Sentence: that seems pretty straightforward
Word: straightforward
# The comment sentence:
{review_sentence}
Output the only comment-related word:
Table 14: Prompts for Extracting Prejudgment Sentences and Words.
cate. subcatrgories Explanation and cases
reasoning question numerical reasoning Questions that require performing basic mathematical operations or solving simple algebraic
equations to arrive at a numerical answer.
Case: What’s two plus three?
code reasoning Questions that require tracing through and executing simple code snippets to determine their
output or behavior when run in a specific programming environment.
Case: What is the output of the following code when run in Python 3 environment: word =
[…]
set reasoning Questions that require applying simple syllogistic reasoning to determine whether elements
belong to sets based on clearly stated relationships.
Case: All squares are quadrilaterals. A shape is a square, is it a quadrilateral?
temporal reasoning Questions that require calculating time durations, ages, or future dates by applying simple
arithmetic operations to temporal information.
Case: How many minutes equal 120 seconds?
spatial reasoning Questions that require determining relative positions, directions, or orientations of objects
in space based on simple spatial relationships.
Case: If a bird is flying above a tree, where is the tree in relation to the bird?
causal reasoning Questions that require determining outcomes by applying simple cause-and-effect
relationships based on given conditional statements.
Case: If ferromagnetic material is placed in a magnetic field, it will become magnetized. An
iron nail was placed next to a strong magnet for some time. Has the nail been magnetized?
natural law reasoning Questions that require applying basic knowledge of physical laws and natural phenomena to
predict simple observable outcomes in everyday scenarios.
Case: Which is faster, an airplane or the propagation of light?
knowledge question geometry facts Questions that require recalling simple and fundamental geometric properties about shapes,
angles, and basic geometric figures.
Case: How many angles does a trapezoid have?
geographic facts Questions that require recalling simple factual information about locations, landmarks,
political divisions, celestial bodies, and other basic geographic knowledge.
Case: Which is the largest continent on Earth?
historical facts Questions that require recalling basic facts about historical events.
Case: Which country first invented paper?
biographical facts Questions that require recalling basic facts about the identities, achievements, and
characteristics of historical figures.
Case: Who proposed the theory of universal gravitation?
measurement units Questions that require recalling simple conversion relationships between
standard units of measurement.
Case: How many centimeters equal 1 meter?
scientific notation Questions that require recalling basic scientific symbols, formulas, and standard units
used in scientific communication.
Case: What is the chemical symbol for oxygen?
creative authorship Questions that require recalling the creators or originators of notable artistic, literary,
musical, and cultural works.
Case: Who is the author of Hamlet?
Table 15: The subcategory descriptions and cases of reasoning questions and knowledge questions.
cate. subcatrgories Explanation and cases
instruction following repetition constraints Questions that require outputting specified characters, words, or phrases a specific number
of times according to simple formatting instructions.
Case: Output the number "7" four times, without using separators.
length constraints Questions that require generating outputs of a specific length or with a specific number
of components based on simple counting constraints.
Case: Output a four-digit number.
character constraints Questions that require generating words or numbers that conform to simple specified
character patterns or formatting rules.
Case: Output a number that begins with 8.
counting constraints Questions that require counting specific characters or elements within
a given text or sequence.
Case: Output the number of letter "y" in the word "yes".
transformation constraints Questions that require modifying text or numbers according to simple formatting or character
substitution rules to produce a transformed output.
Case: Output the word "good" with all letters capitalized directly.
sentence constraints Questions that require generating sentences that conform to simple specified content or
structural requirements.
Case: Give a sentence that contains the phrase "have lunch" directly.
analysis question sentiment classification Questions that require determining whether simple statements express positive or negative
emotions based on the tone and word choice.
Case: Does the sentence "I hate rainy days." express a positive or negative emotion?
named entity recognition Questions that require identifying the correct category of named entities (such as people,
places, organizations, or time expressions) within simple sentences.
Case: In the sentence "Napoleon died in 1821", is "1821" a time or a place name?
language classification Questions that require identifying the language of origin for simple words or phrases
based on their characteristic writing systems or common vocabulary.
Case: Is the word "hello" English or Japanese?
topic classification Questions that require identifying the primary subject matter or thematic category of
simple sentences based on their content and context clues.
Case: Is the topic of the sentence "The stock market rose 2% today" finance or technology?
intent recognition Questions that require determining the communicative purpose behind simple utterances
or statements based on their phrasing and context.
Case: Is the intention of the sentence "I’m sorry I’m late." to apologize or to blame?
syntax classification Questions that require identifying the correct grammatical structure or sentence type of
simple expressions based on their form, punctuation, and communicative function.
Case: Is "Close the door!" an imperative sentence or an interrogative sentence?
grammar classification Questions that require identifying simple grammatical properties (like tense, voice,
or polarity) of sentences based on their structure and verb forms.
Case: Is "The apple was eaten." in active voice or passive voice?
coreference resolution Questions that require identifying which entity a pronoun or reference term refers to in
simple sentences by tracking relationships between words in the text.
Case: In "My computer is broken, and I need to fix it." What does "it" refer to?
Table 16: The subcategory descriptions and cases of instruction following questions and analysis questions.